Two non-mixing liquida having densities p and 2p are put in a container. The height of each liquid is h. A solid cylinder of length l and density d is put in this container. The cylinder floats with its vertical axis and length 0.5 l in the denser liquid. What is density d in terms of p.

Grade:9

1 Answers

Arun
6 years ago
Statement of the question is not clear. I am solving the question based on some considerations.

Please find below the solution to the asked query:

Given that the cylinder is floating in two liquids. Given that the length of the cylinder is L and it is having a part ρL in the denser liquid. So,

ρgAL=AρLndg+A(1−ρ)Ldg⇒ρ=ρnd+(1−ρ)d ⇒ρ=ρnd+d−ρd⇒ρ(1+d−nd)=d ⇒ρ=d(1+d−nd)⇒ρ=d1+d(1−n) 
Hope this information will clear your doubts about the topic.